OpenStudy (anonymous):

Jenny's parents deposit $10,000 for her college tuition in an account with a 3% simple interest rate. How long will it take to earn $3600 in interest?

OpenStudy (texaschic101):

Interest = Principle * rate * time P = 10,000 R = 3% = 0.03 T (we need to solve for t I = 3600 now we sub.. 3600 = (10,000)(0.03)(t) 3600 = 300t -- divide both sides by 300 3600/300 = t 12 = t check.. I = prt 3600 = (10000)(0.03)(12) 3600 = (300)(12) 3600 = 3600 (correct) It would take 12 years